Unemployment: Given the following information, calculate (1) the unemployment rate, (2) the employment rate and (3)...

  1. Unemployment: Given the following information, calculate (1) the unemployment rate, (2) the employment rate and (3) the labor force participation rate.

Total Population = 800,000

Children, Military & Institutionalized = 200,000

Retirees, Stay-at-home parents, students = 100,000

Employed = 475,000

Calculate the Civilian Non-institutionalized Population (CNP):  _____________

Calculate the Civilian Labor Force:  ________________
Answer #1

(a) CNP = Total population - Children, Military & Institutionalized

= 800,000 - 200,000

= 600,000

(b) Civilian labor force (LF) = CNP - Retirees, Stay-at-home parents, students

= 600,000 - 100,000

= 500,000

(c) Unemployment rate = (LF - Employed) / LF

= (500,000 - 475,000) / 500,000

= 25,000 / 500,000

= 0.05

= 5%

(d) Employment rate = 100% - Unemployment rate

= 100% - 5%

= 95%

(e) Labor force participation rate = LF / CNP

= 500,000 / 600,000

= 0.8333

= 83.33%